. The size of the labor force is 400 million. Each month 36% of unemployed workers find jobs and 4% of employed workers lose their jobs.

The number of employed individuals is (in millions)
Answer #1

Answer

Initially Let number of unemployed people be U and number of employed people be E and L be labor force.

Formula E + U = L = 400 million => E + U = 400 million ---------------------------(1)

Now 36% unemployed find jobs and 4% gets unemployed.

(U - (36/100)U) + (E + (4/100)E) = L = 400 million

=> 0.64U + 1.04E = 400 million -------------------------------(2)

From (1) and (2) we get :

E + U = 400 million = 0.64U + 1.04E

=> 0.36U = 0.04E

=> E = 9U

Putting this in (1) we get :

U + 9U = 400 million => U = 40 million => E = 9*40 million = 360 million.

Hence, The number of employed individuals is 360 millions